OpenStudy (anonymous):

the angels of a quadrilateral are in AP whose common difference is 10 degree .find the angles

OpenStudy (mayankdevnani):

let one of the angles be a so other angles will be..... a+10, a+20,a+30 (as the common difference is 10 degree) a + (a+10) + (a+20) + (a+30) = 360 (angle sum property of a quadrilaterral) 4a + 60 = 360 4a = 300 a = 75..................a + 10 = 75 + 10= 85 similarly substitute value of a for all the 4 angles answer will be....... 75,85,95,105 degrees

OpenStudy (mayankdevnani):

ok @mihir786

OpenStudy (mayankdevnani):

@mukushla

OpenStudy (mayankdevnani):

am i right? @mukushla

OpenStudy (anonymous):

yeah :)